# Refined Inference for Asymptotically Linear Estimators with Non-Negligible Second-Order Remainders

### 18 Mar 2026 v2

@@ Abstract
Asymptotically linear estimators in semiparametric models achieve their point-estimation guarantees via a von Mises expansion in which a second-order remainder is declared negligible. Confidence intervals then treat the first-order influence-function term as the sole source of sampling variability. This reasoning is asymptotically exact but can fail materially in finite samples whenever the second-order remainder contributes variation of the same order as the influence-function variance — a regime we call the near-boundary regime, characterized by nuisance estimation operating at or near the product-rate threshold.
We develop a general theory of inference for this regime. Our contributions are: (i) a finite-sample variance decomposition that separates influence-function variance from remainder-induced variance and the covariance between them; (ii) a sandwich consistency theorem that gives a precise necessary and sufficient condition — strong remainder negligibility — for the standard sandwich to be consistent for the total sampling variance, and shows this is strictly stronger than the product-rate condition that guarantees asymptotic linearity; (iii) two refined variance estimators — leave-one-unit-out jackknife and pairs cluster bootstrap — each with full asymptotic validity guarantees in the near-boundary regime, together with a heteroskedasticity-corrected sandwich interpretation that is numerically equivalent to the jackknife Wald interval; and (iv) a clustered-data extension in which the remainder interacts with intra-cluster correlation to produce an analytic formula for sandwich gap amplification.
The theory is developed in a general i.i.d. framework and extended to clustered data. The motivating application is the surrogate-assisted targeted minimum loss estimator (SA-TMLE) for nested bridge functionals under administrative censoring (Li, 2026), where the near-boundary regime arises by design in stepped-wedge cluster-randomized trials. Simulations confirm that the variance ratio $\hat{\rho}=\hat{V}_{\rm JK}/\hat{V}_{\rm Sand}$ is approximately 1.14–1.38 and persistent across all examined cluster counts $J\in\{10,30,50,100\}$, that the sandwich confidence interval undercovers by 4–12 percentage points, and that both refined procedures restore coverage to the nominal level.
Keywords: asymptotic linearity; semiparametric inference; second-order remainder; jackknife; bootstrap; heteroskedasticity correction; targeted learning; nested bridge functional; stepped-wedge trial.
@@ 1 Introduction
1.1 The General Problem
A central achievement of modern semiparametric theory is the asymptotically linear estimator (ALE): an estimator $\hat{\Psi}$ of a functional $\Psi(P_{0})$ of an unknown distribution $P_{0}$ that satisfies
@@ 1.1 The General Problem
where $D^{*}$ is the efficient influence function (EIF) with $\mathbb{E}_{P_{0}}[D^{*}]=0$ and $\mathbb{E}_{P_{0}}[(D^{*})^{2}]=\sigma^{2}_{\mathrm{EIF}}<\infty$. The asymptotic variance of $\hat{\Psi}$ is $\sigma^{2}_{\mathrm{EIF}}/n$, estimated consistently by the empirical variance of the plug-in EIF scores, and Wald intervals achieve nominal coverage in the limit.
This framework encompasses a broad class of practically important estimators: augmented inverse probability weighted (AIPW) estimators (Robins et al., 1994), one-step debiased machine-learning (DML) estimators (Chernozhukov et al., 2018), targeted minimum loss estimators (TMLE) (van der Laan and Rose, 2011), and their clustered-data variants (Balzer et al., 2021). In each case, asymptotic linearity rests on a von Mises expansion argument that controls a second-order remainder $R_{\mathrm{rem}}$ via product- rate conditions on the nuisance estimators.
The inferential logic of (1) is impeccable in the asymptotic limit. But finite-sample confidence interval calibration depends on a condition that is not directly verified: that $R_{\mathrm{rem}}$ is negligible relative to the first-order term not only in probability — as required for consistency — but also in variance, as required for the sandwich to be a valid variance estimator. These are different conditions. Point-estimation consistency requires $R_{\mathrm{rem}}=o_{P}(n^{-1/2})$, which controls the mean of the remainder. Coverage accuracy requires $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$, which controls the second moment. The second condition is strictly stronger than the first and is typically not implied by standard product-rate arguments.
When $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\not\to 0$ — or more concretely, when $\mathrm{Var}(R_{\mathrm{rem}})=O(n^{-1})$ rather than $o(n^{-1})$ — the total sampling variance of $\hat{\Psi}$ contains a non-negligible component beyond $\sigma^{2}_{\mathrm{EIF}}/n$, the sandwich underestimates the true variance, and Wald intervals undercover. We call this the near-boundary regime, since it is characteristic of nuisance estimation at exactly the product-rate boundary.
The problem is not merely theoretical. It has been observed empirically in simulation studies for TMLE (van der Laan and Rose, 2011), doubly robust estimators (Vermeulen and Vansteelandt, 2015), and semiparametric survival estimators (Luedtke and van der Laan, 2017). In each case, the point estimator behaves well while the coverage is below nominal, and the ratio $\widehat{\mathrm{Var}}_{\mathrm{Sand}}/\widehat{\mathrm{Var}}_{\mathrm{emp}}$ is less than one and approximately constant across sample sizes — the signature of near-boundary behaviour identified in Section 3.
1.2 Motivating Application: Surrogate-Assisted Targeted Learning
The concrete motivation for this paper is the surrogate-assisted TMLE (SA-TMLE) of Li (2026) for the average treatment effect under administratively censored delayed outcomes. The estimand is a nested bridge functional identified by integrating an observed-outcome regression over the conditional surrogate distribution:
$\Psi(P_{0})\;=\;\mathbb{E}_{W,t}\Bigl[\mathbb{E}_{S|A=1,W,t}\bigl[\bar{Q}_{Y}(S,1,W,t)\bigr]-\mathbb{E}_{S|A=0,W,t}\bigl[\bar{Q}_{Y}(S,0,W,t)\bigr]\Bigr],$ (2)
where $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ and $\Delta$ is the outcome observation indicator. The SA-TMLE achieves asymptotic linearity under the product-rate condition $\left\|\hat{\bar{Q}}_{Y}-\bar{Q}_{Y}^{0}\right\|_{P_{0}}\cdot\left\|\hat{g}_{\Delta}-g^{0}_{\Delta}\right\|_{P_{0}}=o_{P}(J^{-1/2})$ by absorbing the nested cross-product remainder $R_{SY}$ into the efficient score via a two-stage fluctuation step that circumvents the need to estimate the conditional surrogate density $f_{S}$ directly.
The companion simulation study (Li, 2026, Table 1) reveals a striking pattern. Across $J\in\{10,20,30,50,100\}$ clusters and ICC $\in\{0.01,0.05,0.10,0.20\}$, the SA-TMLE achieves near-zero bias at all configurations, while the sandwich 95% confidence intervals attain only 0.87–0.91 empirical coverage. The ratio of empirical variance to mean sandwich estimate is approximately 1.6 and is constant across all $J$. This pattern cannot be explained by the Berry–Esseen rate $O(J^{-1/2})$ for the first-order CLT, and it does not diminish as $J$ grows to 100: it is a persistent, structural feature of operating at the product-rate boundary with parametric nuisance specifications.
This application has three features that motivate a general theory. First, the second-order remainder $R_{2}$ of the SA-TMLE has a specific bilinear structure — a product of nuisance errors — that is representative of a broad class of semiparametric estimators. Second, the clustered data structure means that the remainder interacts with intra-cluster correlation in an analytically tractable way. Third, the SA-TMLE’s elimination of $R_{SY}$ via the nested fluctuation step means the only remaining remainder is the standard bilinear $R_{2}$, making the problem clean enough to develop a sharp theory.
1.3 Relation to Existing Literature
The phenomenon of non-negligible second-order terms degrading inference has been identified in several distinct literatures, each with a different structural source and remedy.
Kernel-based semiparametric estimators.
Cattaneo and Jansson (2018) study semiparametric estimators whose nonparametric ingredient converges at a slower-than-usual rate due to small bandwidth selection. Their central finding is that the distributional approximation differs from the standard one by a non-negligible bias term, and that some — but not all — bootstrap procedures automatically correct for this bias. The structural parallel to our work is exact: in both settings, the standard first-order approximation fails and resampling provides a remedy. The structural difference is equally exact: in their setting the second-order term inflates the mean of the estimator (bias), while in our near-boundary regime the second-order remainder contributes non-negligible variance while contributing negligible bias. Their nuisance rate is controlled by the analyst via bandwidth choice; ours is dictated by the product-rate constraint inherent to doubly robust semiparametric estimation with nonparametric nuisance. Accordingly, their correction operates through bootstrap bias adjustment, while ours operates through variance estimation.
Finite-dimensional nuisance estimation.
Lok (2024) shows that ignoring finite-dimensional nuisance estimation renders the sandwich conservative (overcoverage), and derives a consistent corrected sandwich. This is structurally opposite to our setting: their nuisance is finite-dimensional and estimated via score equations (leading to underdispersion of the sandwich), while ours is infinite-dimensional and nonparametrically estimated at the product-rate boundary (leading to overdispersion of the total variance relative to the sandwich). Lok explicitly identifies extension to TMLE with Super Learner as future work; our paper addresses precisely this extension in the harder near-boundary regime.
Targeted variance estimation.
Ji et al. (2025) propose a one-step TMLE for the variance of the log causal risk ratio, derived by computing the efficient influence function for the variance functional itself. Their approach is estimator-specific: it requires re-deriving the EIF for each new target parameter and nuisance configuration. Our resampling-based approach (jackknife, bootstrap) applies to any ALE without re-deriving influence functions, and we additionally provide the formal decomposition, the sharp sandwich iff characterization, and the clustered extension. The two approaches are complementary: targeted variance estimation is asymptotically optimal for the specific parameters it targets; our approach trades potential efficiency for generality and transparency.
1.4 Contributions and Outline
This paper makes five contributions.
1. Variance decomposition (Section 3). We derive a finite-sample variance decomposition for general ALEs:
$\mathrm{Var}\bigl(\hat{\Psi}-\Psi(P_{0})\bigr)=\frac{\sigma^{2}_{\mathrm{EIF}}}{n}+\mathrm{Var}(R_{\mathrm{rem}})+2\,\mathrm{Cov}(\bar{D}_{n},R_{\mathrm{rem}})+o_{P}(n^{-1}),$
and characterize the near-boundary regime in terms of the remainder’s $L^{2}$ rate. The constancy of the empirical variance ratio is identified as the finite-sample signature of operating at the product-rate boundary.
2. Sandwich consistency theorem (Section 4). We give a precise necessary and sufficient condition — strong remainder negligibility $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$ — for the sandwich to be consistent for the total variance, and show this is strictly stronger than standard product-rate conditions.
3. Two refined variance estimators (Section 5). We develop:
  1. a.
    The leave-one-unit-out jackknife, whose studentized pivot is self-normalizing: numerator and denominator both track the same total variance $\sigma^{2}_{\mathrm{EIF}}+c_{R}$, so the ratio converges to $N(0,1)$ even though the jackknife estimator itself converges to $\sigma^{2}_{\mathrm{EIF}}/n$.
  2. b.
    The pairs cluster bootstrap, which directly targets the full sampling distribution and achieves second-order accurate coverage via BCa correction.
For each we establish consistency for the total sampling variance and asymptotic validity of the resulting interval. We additionally show that an HC-corrected sandwich — applying the jackknife-to-sandwich ratio as an inflation factor — is numerically equivalent to the jackknife Wald interval, providing a useful computational interpretation.
4. Clustered extension (Section 6). We show that in clustered data the sandwich gap is amplified by a factor proportional to the ICC and the cluster size, providing the first analytic explanation for why the near-boundary regime is especially consequential in SW-CRTs.
5. Simulation study (Section LABEL:sec:simulation). We provide simulation evidence for two applications: the SA-TMLE in SW-CRTs and a cross-fitted AIPW estimator. Both demonstrate the near-boundary constant variance ratio, the failure of the sandwich, and the recovery of correct coverage by the jackknife and bootstrap. The variance ratio diagnostic $\hat{\rho}=\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Sand}}$ is documented as a practical tool for detecting near-boundary operation in applied work.
Section 2 establishes the general framework. The paper is written for a reader familiar with semiparametric efficiency theory (Bickel et al., 1993; van der Vaart, 1998) and targeted learning (van der Laan and Rose, 2011).
@@ 5.1.2 Leave-One-Out Expansion
Sufficient condition for (S). Condition (S) holds whenever the nuisance estimator $\hat{\eta}_{k}$ is a leave-one-out stable empirical risk minimizer: that is, when $\hat{\eta}_{k}$ solves $\min_{\eta}\mathbb{P}_{n}\ell(\eta;O)$ over a function class $\mathcal{F}$ and the loss $\ell$ satisfies a uniform stability bound $\sup_{\eta}|(\mathbb{P}_{n}-\mathbb{P}_{n}^{(-i)})\nabla_{\eta}\ell(\eta;\cdot)|=O_{P}(n^{-1/2})$ (cf. Bousquet and Elisseeff, 2002, Chapter 3). In the SA-TMLE context, cross-fitted Super Learner satisfies this by construction: removing one observation from a fold changes the fold-specific nuisance by at most $O_{P}(n_{\mathrm{fold}}^{-1/2})=O_{P}(n^{-1/2})$ in $L^{2}$ norm under Donsker or entropy conditions on the library.
@@
Step 1: influence-function term. Write $\hat{\Psi}^{(-i)}-\Psi(P_{0})=(n-1)^{-1}\sum_{k\neq i}D^{*}(O_{k})+R_{\mathrm{rem}}^{(-i)}$ by asymptotic linearity of the LOO estimator, and $\hat{\Psi}-\Psi(P_{0})=n^{-1}\sum_{k}D^{*}(O_{k})+R_{\mathrm{rem}}$. Subtracting:
$\hat{\Psi}^{(-i)}-\hat{\Psi}$ $=\Bigl(\frac{1}{n-1}\sum_{k\neq i}-\frac{1}{n}\sum_{k}\Bigr)D^{*}(O_{k})+\bigl(R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}\bigr)$
$=-\frac{1}{n}D^{*}(O_{i})+\underbrace{\frac{1}{n(n-1)}\sum_{k\neq i}D^{*}(O_{k})}_{\displaystyle=:\,a_{i}}+\underbrace{\bigl(R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}\bigr)}_{\displaystyle=:\,b_{i}}.$ (10)
Step 2: rate of $a_{i}$. Since $\|D^{*}\|_{\infty}\leq M$ (Assumption 2.3), $|a_{i}|\leq M/n=O(n^{-1})$ deterministically, so $\max_{i}|a_{i}|=O(n^{-1})$.
Step 3: rate of $b_{i}$ via stability condition (S). By the bilinear structure of $R_{\mathrm{rem}}$ (Definition 2.1):
$b_{i}$ $=R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}=\int\bigl[(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})(\hat{\eta}_{2}^{(-i)}-\eta_{2}^{0})-(\hat{\eta}_{1}-\eta_{1}^{0})(\hat{\eta}_{2}-\eta_{2}^{0})\bigr]\,dP_{0}+o_{P}(r_{n}).$
Adding and subtracting $(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})(\hat{\eta}_{2}-\eta_{2}^{0})$:
$b_{i}$ $=\int(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})\bigl[(\hat{\eta}_{2}^{(-i)}-\eta_{2}^{0})-(\hat{\eta}_{2}-\eta_{2}^{0})\bigr]\,dP_{0}+\int\bigl[(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})-(\hat{\eta}_{1}-\eta_{1}^{0})\bigr](\hat{\eta}_{2}-\eta_{2}^{0})\,dP_{0}+o_{P}(r_{n}).$
By Cauchy–Schwarz, each integral is bounded by $\|\hat{\eta}_{k}-\eta_{k}^{0}\|_{P_{0}}\cdot\|\hat{\eta}_{k}^{(-i)}-\hat{\eta}_{k}\|_{P_{0}}=O_{P}(r_{n})\cdot O_{P}(n^{-1/2})=O_{P}(r_{n}/n^{1/2})$, using Assumption 2.4 and condition (S). At the near-boundary, $r_{n}=O(n^{-1/2})$, so $b_{i}=O_{P}(n^{-1})$ uniformly in $i$.
Step 4: quadratic sum. Set $\delta_{i}=a_{i}+b_{i}$. Since $|a_{i}|=O(n^{-1})$ deterministically and $|b_{i}|=O_{P}(n^{-1})$ uniformly in $i$,
$\max_{i}|\delta_{i}|\;\leq\;\max_{i}|a_{i}|+\max_{i}|b_{i}|\;=\;O(n^{-1})+O_{P}(n^{-1})=O_{P}(n^{-1}).$
For the quadratic sum, since $|b_{i}|\leq\|\hat{\eta}_{1}-\eta_{1}^{0}\|_{P_{0}}\|\hat{\eta}_{2}^{(-i)}-\hat{\eta}_{2}\|_{P_{0}}+\|\hat{\eta}_{1}^{(-i)}-\hat{\eta}_{1}\|_{P_{0}}\|\hat{\eta}_{2}-\eta_{2}^{0}\|_{P_{0}}+o_{P}(r_{n})=O_{P}(r_{n}/n^{1/2})$ uniformly, we have $b_{i}^{2}=O_{P}(r_{n}^{2}/n)$ uniformly, so $n^{-1}\sum_{i}b_{i}^{2}=O_{P}(r_{n}^{2}/n)$, and the stated rate follows. ∎

### 23 Mar 2026 v3

@@ Abstract
Asymptotically linear estimators in semiparametric models use a von Mises expansion to reduce inference to the influence-function variance. This reduction is valid when the second-order remainder is negligible in variance, a condition that is not implied by the product-rate requirement guaranteeing asymptotic linearity. When the remainder contributes non-negligible variance, the standard sandwich underestimates the total sampling variance and Wald intervals undercover; we call this the near-boundary regime. We derive a finite-sample variance decomposition separating influence-function and remainder components, characterize the sandwich consistency condition sharply, and establish asymptotic validity of the leave-one-out jackknife and pairs cluster bootstrap in this regime. Jackknife validity follows from a self-normalization argument; bootstrap validity is established directly. An analytic formula for the amplification of the sandwich gap by intra-cluster correlation is derived for clustered data. A simulation study using a surrogate-assisted targeted learning estimator in stepped-wedge cluster-randomized trials illustrates the regime: the variance ratio $\hat{V}_{\rm JK}/\hat{V}_{\rm Sand}$ is 1.14–1.38 and persistent across cluster counts, and the refined procedures substantially improve coverage.
Keywords: asymptotic linearity; semiparametric inference; second-order remainder; jackknife; bootstrap; variance decomposition; stepped-wedge trial.
@@ 1 Introduction
Asymptotically linear estimators (ALEs) in semiparametric models satisfy
@@
where $D^{*}$ is the efficient influence function (EIF) with variance $\sigma^{2}_{\mathrm{EIF}}$. This framework covers AIPW (Robins et al., 1994), DML (Chernozhukov et al., 2018), and TMLE (van der Laan and Rose, 2011). Asymptotic linearity rests on a von Mises expansion in which a second-order remainder $R_{\mathrm{rem}}$ is controlled via product-rate conditions on nuisance estimators. The standard Wald interval then uses the sandwich variance $\widehat{\sigma}^{2}_{\mathrm{EIF}}/n$ as an estimate of the total sampling variance.
This reasoning is asymptotically exact, but finite-sample coverage depends on a condition that is not implied by asymptotic linearity itself: the second-moment negligibility $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$. Standard product-rate conditions guarantee $R_{\mathrm{rem}}=o_{P}(n^{-1/2})$, which suffices for consistency of the point estimator. First-order variance estimation requires additionally that $\mathrm{Var}(R_{\mathrm{rem}})=o(n^{-1})$, a stronger condition that fails when nuisance estimation operates at the product-rate boundary. When $\mathrm{Var}(R_{\mathrm{rem}})=O(n^{-1})$, the total sampling variance of $\hat{\Psi}$ contains a structural component $c_{R}/n>0$ beyond $\sigma^{2}_{\mathrm{EIF}}/n$, the sandwich underestimates the total variance, and Wald intervals undercover. We call this the near-boundary regime.
The gap between asymptotic linearity and valid first-order inference has been noted empirically for TMLE (van der Laan and Rose, 2011), doubly robust estimators (Vermeulen and Vansteelandt, 2015), and semiparametric survival estimators (Luedtke and van der Laan, 2017), but has not been given a general theoretical treatment. matsushita2020jackknife identify related nonstandard behaviour for jackknife empirical likelihood in the kernel semiparametric setting; their structural problem is different (nuisance rate controlled via bandwidth), but the remedy—resampling—is analogous. Cattaneo and Jansson (2018) study non-negligible bias from slow nuisance rates; our near-boundary remainder contributes non-negligible variance while being asymptotically unbiased, a structurally distinct problem.
The motivating application is the surrogate-assisted TMLE (SA-TMLE) of Author (2026) for nested bridge functionals under administrative censoring in stepped-wedge cluster-randomized trials. Across $J\in\{10,\ldots,100\}$ clusters and ICC $\in\{0.01,\ldots,0.20\}$, the SA-TMLE achieves near-zero bias while the sandwich 95% intervals attain only 0.87–0.91 empirical coverage, with the variance ratio $\hat{\rho}=\widehat{\mathrm{Var}}_{\mathrm{JK}}/\widehat{\mathrm{Var}}_{\mathrm{Sand}}\approx 1.6$ constant across all $J$ — a structural signature of the near-boundary regime.
Contributions.
We develop a sharp general framework for inference in the near-boundary regime. (i) A finite-sample variance decomposition separates influence-function variance from remainder-induced variance and their covariance. (ii) A sandwich consistency characterization (Theorem 4.1) gives a necessary and sufficient condition for the sandwich to estimate total variance: $n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to 0$. (iii) Two refined variance estimators — leave-one-unit-out jackknife and pairs cluster bootstrap — remain valid in the near-boundary regime; bootstrap validity is established directly, jackknife validity via a self-normalization argument in which numerator and denominator both track $\sigma^{2}_{\mathrm{EIF}}+c_{R}$. An HC-corrected sandwich is shown to be numerically equivalent to the jackknife Wald interval. (iv) A clustered-data extension derives an analytic formula for ICC amplification of the sandwich gap.
Section 2 establishes notation; Sections 34 give the decomposition and sandwich characterization; Section 5 develops the refined estimators; Section 6 treats clustered data; Section 7 presents simulations; Section 8 concludes.
@@ 5.1.2 Leave-One-Out Expansion
Sufficient condition for (S). A sufficient route to Condition (S) is leave-one-out stability of an empirical risk minimizer: when $\hat{\eta}_{k}$ solves $\min_{\eta}\mathbb{P}_{n}\ell(\eta;O)$ over a function class $\mathcal{F}$ and the loss satisfies a uniform gradient stability bound $\sup_{\eta}|(\mathbb{P}_{n}-\mathbb{P}_{n}^{(-i)})\nabla_{\eta}\ell(\eta;\cdot)|=O_{P}(n^{-1/2})$ (cf. Bousquet and Elisseeff, 2002, Chapter 3), then (S) holds under standard regularity. In the SA-TMLE context, cross-fitted Super Learner may satisfy this under additional entropy and stability conditions on the library; general verification is case-specific.
@@
Subtracting the ALE expansions of $\hat{\Psi}$ and $\hat{\Psi}^{(-i)}$:
$\hat{\Psi}^{(-i)}-\hat{\Psi}=-\frac{1}{n}D^{*}(O_{i})+\underbrace{\frac{1}{n(n-1)}\sum_{k\neq i}D^{*}(O_{k})}_{a_{i}}+\underbrace{R_{\mathrm{rem}}^{(-i)}-R_{\mathrm{rem}}}_{b_{i}}.$
Since $\|D^{*}\|_{\infty}\leq M$, $|a_{i}|\leq M/n$ deterministically. For $b_{i}$: the bilinear structure of $R_{\mathrm{rem}}$ gives $b_{i}=\int(\hat{\eta}_{1}^{(-i)}-\eta_{1}^{0})(\hat{\eta}_{2}^{(-i)}-\hat{\eta}_{2})\,dP_{0}+\int(\hat{\eta}_{1}^{(-i)}-\hat{\eta}_{1})(\hat{\eta}_{2}-\eta_{2}^{0})\,dP_{0}+o_{P}(r_{n})$. By Cauchy–Schwarz and condition (S), each term is $O_{P}(r_{n}\cdot n^{-1/2})$, giving $\max_{i}|b_{i}|=O_{P}(r_{n}/n^{1/2})$. Setting $\delta_{i}=a_{i}+b_{i}$ gives $\max_{i}|\delta_{i}|=O_{P}(n^{-1})$ at $r_{n}=O(n^{-1/2})$, and $n^{-1}\sum_{i}\delta_{i}^{2}=O_{P}(r_{n}^{2}/n)$. ∎